Find sin 0 if cos 0= 24/25
Answer by amarjeeth123(574)

cos theta=24/25
We have sin^2theta+cos^2theta=1
sin theta=√1-cos^2theta=√1-(24/25)^2=√1-576/625=√49/625=7/25
Answer:sin theta=7/25